OpenStudy (anonymous):

use the diagonals to determine whether a parallelogram with the given vertices is a rectangle, rhombus, or square. Give all the names that aply.. J(-9, -7) K(-4,-2) L(3,-3) M(-2,-8)

OpenStudy (anonymous):



OpenStudy (anonymous):

no no that's all wrong

OpenStudy (anonymous):

i'm serious guy's..i need help..stephen@west can you help? or someone please?

OpenStudy (anonymous):

yes

OpenStudy (accessdenied):

Graph the four vertices. Check: the length of both diagonals and the slopes of both diagonals. A rectangle has congruent diagonals. A rhombus has perpendicular diagonals. A square has properties of both a rectangle and a rhombus, so it would have both congruent and perpendicular diagonals.

OpenStudy (anonymous):

It's a quadralateral

OpenStudy (anonymous):

AccessDenied..isn't it a rhombus?

OpenStudy (accessdenied):

I haven't checked yet -- I'll go figure it out.

OpenStudy (anonymous):

ok

OpenStudy (accessdenied):

Yes, it is a rhombus.

OpenStudy (anonymous):

the slope of J and L = -2/3 and the slope of K and M are -1/2?

OpenStudy (accessdenied):

J(-9, -7) K(-4,-2) L(3,-3) M(-2,-8) m(JL) = (y2 - y1) / (x2 - x1) = (-3 - (-7)) / (3 - (-9)) = (-3 + 7) / (3 + 9) = 4 / 12 = 1 / 3 m(KM) = (y2 - y1)/(x2 - x1) = (-8 - (-2)) / (-2 - (-4)) = (-8 + 2) / (-2 + 4) = (-6)/(2) = -3/1 = -3 m(KM) * m(KL) = (1/3) * (-3) = -1 (perpendicular)

OpenStudy (anonymous):

Thank's! I don't know how i got the slope wrong I wrote it right but got the wrong answer.. :O ..i gett it know! Thank you very much!!! (:

OpenStudy (accessdenied):

No problem! I should note that I made a small typo on the last line -- it should say, "m(JL)", and not "m(KL)".

OpenStudy (anonymous):

Ok thank's!